# Lennart Berg and Torben Wiedemann

### v1

@@ 4 Foldings
If $\bar{\pi}$ is linear or (induced by) the scaling map $\phi\colon v\mapsto\frac{1}{\lVert v\rVert}v$, then the condition in 4.16 that $\bar{\pi}(\mathord{]A,B[})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}$ (or $\bar{\pi}(\mathord{]A,B[}_{\text{cry}})\subseteq\mathord{]\bar{\pi}(A),\bar{\pi}(B)[}_{\text{cry}}$) is satisfied.
Proposition 4.18.
The group $\operatorname{Ch}(D_{6},\mathcal{R})$ has crystallographic $GH_{3}$-commutator relations with root groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$, and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ is a (non-crystallographic) $H_{3}$-grading of $\operatorname{Ch}(D_{6},\mathcal{R})$.
Proof.
The groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ satisfy Axiom 3.3 (ii) by 4.7 and because all groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ are abelian. Further, by 4.16 and 4.17, $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ and $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ also satisfy Axiom 3.3 (i) and $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ even satisfies the crystallographic version of this axiom.
By 4.14, Weyl elements in $(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}}$ exist. Lastly, note that the positive systems $\Pi_{D}$ and $\Pi_{H}$ in 4.8 satisfy $\tilde{U}_{\Pi_{H}}^{H}\cap\tilde{U}_{-\Pi_{H}}^{H}=\tilde{U}_{\Pi_{D}}^{D}\cap\tilde{U}_{-\Pi_{D}}^{D}=\{1\}$ because $(\tilde{U}_{\beta}^{D})_{\beta\in D_{6}}$ is the grading of a Chevalley group (see [Ste16, Corollary 3 of Lemma 18]). By 3.6, this finishes the proof. ∎
Remark 4.19.
The groups $(\tilde{U}_{\beta}^{GH})_{\beta\in GH_{3}}$ satisfy $\tilde{U}_{\Pi_{GH}}^{GH}\cap\tilde{U}_{-\Pi_{GH}}^{GH}=\{1\}$ as well, but one can show that they do not have Weyl elements.
Now that we have an explicit example of an $H_{3}$-graded group at hand, we can compute the commutator formulas in this group.
Notation 4.20.
For all non-proportional roots $\alpha,\gamma\in H_{3}$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we define a map $\tilde{\psi}_{\alpha,\gamma}^{\beta}\colon(\mathcal{R}\times\mathcal{R})^{2}\rightarrow\mathcal{R}\times\mathcal{R}$ by the property that
$[\tilde{\theta}_{\alpha}^{H}(x_{1},y_{1}),\tilde{\theta}_{\gamma}^{H}(x_{2},y_{2})]_{\beta}=\tilde{\theta}_{\beta}^{H}\big{(}\tilde{\psi}_{\alpha,\gamma}^{\beta}\big{(}(x_{1},y_{1}),(x_{2},y_{2})\big{)}\big{)}$
for all $x_{1},x_{2},y_{1},y_{2}\in\mathcal{R}$. If $\mathord{]\alpha,\gamma[}$ has only one element $\beta$, we will write $\tilde{\psi}_{\alpha,\gamma}$ in place of $\tilde{\psi}_{\alpha,\gamma}^{\beta}$.
Remark 4.21.
Let $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}\in\Delta_{H}$ and denote the corresponding $H_{2}$-quintuple by $(\alpha,\beta,\gamma,\delta,\varepsilon)$. We can determine the maps $\tilde{\psi}_{\zeta,\xi}^{\rho}$ for $\zeta,\xi\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ and $\rho\in\mathord{]\zeta,\xi[}$ by straightforward matrix computations (see 2.11). The results are given in Figure 4.
[…]
Figure 4: The commutator relations in $\operatorname{Ch}(D_{6},R)$ for the $H_{3}$-grading $(\tilde{U}_{\zeta}^{H})_{\zeta\in H_{3}}$. See also 4.21.
Remark 4.22.
The family $(\tilde{U}_{\xi}^{GH})_{\xi\in GH_{3}}$ should be regarded as a refinement of the $H_{3}$-grading $(\tilde{U}_{\zeta}^{H})_{\zeta\in H_{3}}$ because its commutator relations are crystallographic. For example, let $(\alpha,\beta,\gamma,\delta,\varepsilon)$ be an $H_{2}$-quintuple in $GH_{3}$ and put $\xi^{\prime}\coloneqq\phi(\xi)$ for all $\xi\in GH_{3}$. Then $\tilde{U}_{\alpha^{\prime}}^{H}=\tilde{U}_{\alpha}^{GH}\tilde{U}_{\tau\alpha}^{GH}$ and $\tilde{U}_{\delta^{\prime}}^{H}=\tilde{U}_{\delta}^{GH}\tilde{U}_{\tau\delta}^{GH}$, so arbitrary $y_{\alpha^{\prime}}\in\tilde{U}_{\alpha^{\prime}}^{H}$ and $y_{\delta^{\prime}}\in\tilde{U}_{\delta^{\prime}}^{H}$ can be written as $y_{\alpha^{\prime}}=x_{\alpha}x_{\tau\alpha}$ and $y_{\delta^{\prime}}=x_{\delta}x_{\tau\delta}$ where $x_{\xi}\in\tilde{U}_{\xi}^{GH}$ for all $\xi\in\{\alpha,\tau\alpha,\delta,\tau\delta\}$. By relations 2.2 (v) and 2.2 (vi),
$[y_{\alpha^{\prime}},y_{\delta^{\prime}}]$ $=[x_{\alpha}x_{\tau\alpha},x_{\delta}x_{\tau\delta}]=[x_{\alpha},x_{\delta}x_{\tau\delta}]^{x_{\tau\delta}}[x_{\tau\alpha},x_{\delta}x_{\tau\delta}]$
Using the crystallographic $GH_{3}$-commutator relations and the crystallographic adjacency relations that we can read off from Figure 3, we infer that
$[y_{\alpha^{\prime}},y_{\delta^{\prime}}]=[x_{\alpha},x_{\tau\delta}][x_{\tau\alpha},x_{\delta}]\in\tilde{U}_{\tau\gamma}^{GH}\tilde{U}_{\tau\beta}^{GH}.$
We conclude that $[\tilde{U}_{\alpha^{\prime}}^{H},\tilde{U}_{\delta^{\prime}}^{H}]\subseteq\tilde{U}_{\tau\gamma}^{GH}\tilde{U}_{\tau\beta}^{GH}$, which is a refinement of the non-crystallographic commutator relation
$[\tilde{U}_{\alpha^{\prime}}^{H},\tilde{U}_{\delta^{\prime}}^{H}]\subseteq\tilde{U}_{\beta}^{H}\tilde{U}_{\gamma}^{H}=\tilde{U}_{\beta}^{GH}\tilde{U}_{\tau\beta}^{GH}\tilde{U}_{\gamma}^{GH}\tilde{U}_{\tau\gamma}^{GH}.$
Remark 4.23.
In a similar way as in this section, one can fold a Chevalley group of type $E_{8}$ along the diagram in Figure 1(c) to obtain an $H_{4}$-graded group.
5 The action of Weyl elements in $H_{3}$-graded groups
@@
The goal of this section is 5.20, which provides formulas for the actions of squares of Weyl elements on all root groups in $G$. Namely, we will see that these actions “factor through” an action of the group $A\coloneqq\{\pm 1\}^{2}$, the so-called twisting group of $G$ (see 6.11). The existence of a twisting group is the main prerequisite that is necessary to apply the parametrisation theorem in Section 6.
@@
Throughout this section, we will often implicitly use 3.7 to reduce statements about $H_{3}$-graded groups to statements about its rank-2 subgroups. Similarly, we will see in 5.22 that all results in this section transfer to the case of $H_{4}$-graded groups without difficulty. Note that all rank-2 subsystem of $H_{3}$ are of type $H_{2}$, $A_{2}$ or $A_{1}\times A_{1}$. In particular, pairs of orthogonal roots are necessarily adjacent. Further, we will frequently apply many statements in this section to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$ in place of $(\alpha,\beta,\gamma,\delta,\epsilon)$, which is legitimate because $(\alpha,\beta,\gamma,\delta,\epsilon)$ is chosen arbitrarily.
The following results summarise what is known from the case of $A_{2}$-gradings.
@@
    The root group $U_{\zeta}$ is abelian even if we do not explicitly require Axiom 3.3 (ii).
@@
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Let $\xi\in H_{3}$ such that $(\zeta,\xi)$ is an $A_{2}$-pair and let $x_{\xi}\in U_{\xi}$. Then $x_{\xi}^{w_{\zeta}}=[x_{\xi},b_{\zeta}]$.
@@
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Then $(a_{-\zeta},b_{\zeta},a_{-\zeta})$ is balanced in the sense of [Wie23, 2.2.14]. That is, we have $a_{-\zeta}^{w_{\zeta}}=b_{\zeta}$, $b_{\zeta}^{w_{\zeta}}=a_{-\zeta}$ and $w_{\zeta}=b_{\zeta}a_{-\zeta}b_{\zeta}$.
@@
Since $\alpha$ is contained in an $A_{2}$-subsystem by 2.12, this follows from the corresponding assertions about $A_{2}$-graded groups in [Wie23, 5.4.3, 5.4.9, 5.4.10]. For the first assertion, see also 3.5. ∎
@@
If $\Psi_{\zeta,\xi}$ is of type $A_{1}\times A_{1}$, then $U_{\xi}$ is centralised by $U_{\zeta}$ and $U_{-\zeta}$ and hence by $w$. The remaining cases are covered by [Wie23, 5.4.13, 5.4.15]. ∎
@@
for some $z_{\gamma}\in U_{\gamma}$ and $z_{\delta}\in U_{\delta}$. By 3.10, the first equation in i follows. Since $U_{\beta}$ is abelian, the second equation in i holds as well. Finally, assertion ii follows from i (applied to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$) and 3.13:
@@
This is a similar computation as in 5.5. See [Wie23, 7.4.4] for details. ∎
@@
For (ii), recall from 3.9 that $w_{\alpha}$ is also a ($-\alpha$)-Weyl element with factorisation $w_{\alpha}=b_{\alpha}a_{-\alpha}a_{-\alpha}^{w_{\alpha}}$. Hence by two applications of (i),
@@
Using the factorisation $w_{\alpha}^{-1}=(a_{-\alpha}^{-1})^{w_{\alpha}}a_{-\alpha}^{-1}b_{\alpha}^{-1}$ of $w_{\alpha}^{-1}$ as a $(-\alpha)$-Weyl element (see 3.9), an application of (i) to the $H_{2}$-quintuple $(-\alpha,\varepsilon,\delta,\gamma,\beta)$ yields $x_{\varepsilon}=\left[a_{-\alpha}^{-1},[b_{\alpha},x_{\varepsilon}]_{\beta}\right]_{\varepsilon}$. This finishes the proof. ∎
@@
Since $w_{\alpha}$ is also a $(-\alpha)$-Weyl element by 3.9 (iii), we may assume that $\zeta\in\{\beta,\gamma,\delta,\varepsilon\}$. In each case, we will derive a formula for the action of $w_{\alpha}^{2}$ on $U_{\zeta}$ which does not depend on the choice of $w_{\alpha}$, which proves the first assertion. It then follows that $w_{\alpha}^{2}$ and $(w_{\alpha}^{-1})^{2}$ act identically on $U_{\zeta}$, which says precisely that $w_{\alpha}^{4}$ acts trivially on $U_{\zeta}$.
Recall from 2.12 that $\zeta$ lies in a unique $H_{2}$-subsystem $\tilde{\Psi}$ different from $\Psi$. We choose an $H_{2}$-quintuple $(\tilde{\alpha},\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in $\tilde{\Psi}$ such that $\zeta=\tilde{\beta}$. Since any rank-2 subsystem is uniquely determined by two non-proportional roots in this subsystem, we have $\Psi\cap\tilde{\Psi}=\{\zeta,-\zeta\}=\{\tilde{\beta},-\tilde{\beta}\}$ and
@@
Now let $x_{\zeta}\in U_{\zeta}$ and put $w\coloneqq w_{\alpha}^{2}$. By 5.8 we can write $x_{\zeta}=[x_{\tilde{\alpha}},x_{\tilde{\varepsilon}}]_{\zeta}$ for some $x_{\tilde{\alpha}}\in U_{\tilde{\alpha}}$ and $x_{\tilde{\varepsilon}}\in U_{\tilde{\varepsilon}}$. Hence by 3.13,
@@
If $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$, then 5.4 provides formulas for the action of $w$ on $U_{\tilde{\alpha}}$ and $U_{\tilde{\varepsilon}}$ which do not depend on the choice of $w_{\alpha}$. Thus the first assertion holds in this case. Further, since $\alpha$ lies in exactly two $H_{2}$-subsystems by 2.12, it follows from (5.10) that at most one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ can be of type $H_{2}$.
We remain with the case that one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ is of type $H_{2}$. Again by applying 5.8, but to the $H_{2}$-pair $(\tilde{\gamma},-\tilde{\delta})$, we can write $x_{\zeta}=[x_{\tilde{\gamma}},x_{-\tilde{\delta}}]_{\zeta}$ for some $x_{\tilde{\gamma}}\in U_{\tilde{\gamma}}$ and $x_{-\tilde{\delta}}\in U_{-\tilde{\delta}}$. As above, we then have $x_{\zeta}^{w}=[x_{\tilde{\gamma}}^{w},x_{-\tilde{\delta}}^{w}]_{\zeta}$. By our case assumption, 2.12 and (5.10), both $\Psi_{\alpha,\tilde{\gamma}}$ and $\Psi_{\alpha,-\tilde{\delta}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$. By 5.4, we conclude that the first assertion holds in this case as well, which finishes the proof. ∎
@@
By 2.12 iii, there are exactly two positive roots in $H_{3}\cap\alpha^{\perp}$, so the uniqueness assertion is trivial. Let $\rho$ denote any root in $H_{3}\cap\alpha^{\perp}$. Then $\rho\notin\Psi$ and hence $\rho\notin\langle\Psi\rangle_{\mathbb{R}}$. By the defining formula for reflections, it follows that for all $\xi\in\Psi$, we have either $\xi^{\sigma(\rho)}=\xi$ or $\xi^{\sigma(\rho)}\notin\Psi$. At the same time, by 2.12 iii, $\sigma_{\rho}$ cannot fix all roots in $\Psi$. Since $\alpha^{\sigma(\rho)}=\alpha$, we conclude that $\Psi^{\sigma(\rho)}$ is the unique $H_{2}$-subsystem of $H_{3}$ containing $\alpha$ which is different from $\Psi$. That is, $\Psi^{\sigma(\rho)}=\tilde{\Psi}$. Since $\sigma_{\rho}$ is orthogonal and $\alpha^{\sigma(\rho)}=\alpha$, we must have $\beta^{\sigma(\rho)}\in\{\tilde{\beta},-\tilde{\varepsilon}\}$.
@@
Let $\eta,\vartheta\in H_{3}\in\alpha^{\perp}$ be as in 5.14 and choose Weyl elements $w_{\eta}\in M_{\eta}$ and $w_{\vartheta}\in M_{\vartheta}$. At first, we consider the case that $\xi$ and $\zeta$ lie in a common $H_{2}$-subsystem. Without loss of generality, we assume that both lie in $\Psi$. Further, by 3.9 (iii) and 5.9, we may assume that $\zeta\neq\xi$ and that both $\xi,\zeta$ are positive roots. Thus either $\{\xi,\zeta\}=\{\beta,\varepsilon\}$ or $\{\xi,\zeta\}=\{\gamma,\delta\}$. In any case, $\xi^{\sigma(\eta)\sigma(\vartheta)}=-\zeta$ by 5.14, so it follows from 3.9 (ii) that
@@
Note that if $\rho$ has the desired properties, then $-\rho$ has the desired properties for $(\alpha,\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in 5.13 replaced by $(\alpha,-\tilde{\varepsilon},-\tilde{\delta},-\tilde{\gamma},-\tilde{\beta})$. Thus for each fixed root $\alpha$, there are only four cases to consider: Namely, there are two ways to choose $\Psi$ and two ways to choose an $H_{2}$-quintuple starting from $\alpha$ in $\Psi$. It is a straightforward computation (see 2.11) to prove the assertion in each of these cases for one fixed choice of $\alpha$. Then it follows from 2.9 that it holds for all possible choices of $\alpha$. ∎
@@
Let $x_{\alpha}\in U_{\alpha}$ be arbitrary. Using 5.8, we can write $x_{\alpha}=[x_{\beta},x_{-\gamma}]_{\alpha}$ for some $x_{\beta}\in U_{\beta},~{}x_{-\gamma}\in U_{-\gamma}$. Choose $a_{-\beta}\in U_{-\beta}$ and $b_{\beta}\in U_{\beta}$ such that $w_{\beta}=a_{-\beta}b_{\beta}a_{-\beta}$. Applying 5.7 (ii) to the $(-\beta)$-Weyl element $w_{\beta}=b_{\beta}a_{-\beta}b_{\beta}^{w_{\beta}}$ from 3.9 (iii), we see that

### v2

@@ 4 Foldings
For any root system $\Phi$, every Chevalley group $(G,(U_{\alpha})_{\alpha\in\Phi})$ satisfies that $U_{\Pi}\cap U_{-\Pi}=\{1_{G}\}$. Further, every $X$-graded group with this property satisfies the additional assumption on $U_{\Pi}^{X}\cap U_{-\alpha_{1}}^{X}U_{-\alpha_{2}}^{X}$ in 4.32. In fact, we are not aware of any examples of root graded groups which do not satisfy $U_{\Pi}\cap U_{-\Pi}=\{1_{G}\}$.
5 The action of Weyl elements in $H_{3}$-graded
groups
@@ 5 The action of Weyl elements in $H_{3}$-graded
The goal of this section is 5.20, which provides formulas for the actions of squares of Weyl elements on all root groups in $G$. Namely, we will see that these actions “factor through” an action of the group $A\coloneqq\{\pm 1\}^{2}$, the so-called twisting group of $G$ (see 6.13). The existence of a twisting group is the main prerequisite that is necessary to apply the parametrisation theorem in Section 6.
@@
Throughout this section, we will often implicitly use 3.8 to reduce statements about $H_{3}$-graded groups to statements about its rank-2 subgroups. Similarly, we will see in 5.21 that all results in this section transfer to the case of $H_{4}$-graded groups without difficulty. Note that all rank-2 subsystem of $H_{3}$ are of type $H_{2}$, $A_{2}$ or $A_{1}\times A_{1}$. In particular, pairs of orthogonal roots are necessarily adjacent. Further, we will frequently apply many statements in this section to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$ in place of $(\alpha,\beta,\gamma,\delta,\epsilon)$, which is legitimate because $(\alpha,\beta,\gamma,\delta,\epsilon)$ is chosen arbitrarily.
The following result summarises what is known from the case of $A_{2}$-gradings.
@@
    The root group $U_{\zeta}$ is abelian.
@@
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Let $\xi\in H_{3}$ such that $(\zeta,\xi)$ is an $A_{2}$-pair and let $x_{\xi}\in U_{\xi}$. Then $x_{\xi}^{w_{\zeta}}=[x_{\xi},b_{\zeta}]$ and $x_{\xi}=[[x_{\xi},b_{\zeta}],a_{-\zeta}]^{-1}$. In particular, $U_{\xi+\zeta}=U_{\xi}^{w_{\zeta}}=[U_{\xi},U_{\zeta}]$.
@@
    Let $w_{\zeta}=a_{-\zeta}b_{\zeta}a_{-\zeta}$ be a $\zeta$-Weyl element with $a_{-\zeta}\in U_{-\zeta}$ and $b_{\zeta}\in U_{\zeta}$. Then $(a_{-\zeta},b_{\zeta},a_{-\zeta})$ is balanced in the sense of [Wie24, 2.2.14]. That is, we have $a_{-\zeta}^{w_{\zeta}}=b_{\zeta}$, $b_{\zeta}^{w_{\zeta}}=a_{-\zeta}$ and $w_{\zeta}=b_{\zeta}a_{-\zeta}b_{\zeta}$.
@@
Since $\alpha$ lies in an $A_{2}$-subsystem by 2.19, this follows from the corresponding assertions about $A_{2}$-graded groups in [Wie24, 5.4.3, 5.4.9, 5.4.10]. ∎
@@
If $\Psi_{\zeta,\xi}$ is of type $A_{1}\times A_{1}$, then $U_{\xi}$ is centralised by $U_{\zeta}$ and $U_{-\zeta}$ and hence by $w$. The remaining cases are covered by [Wie24, 5.4.13, 5.4.15]. ∎
@@
for some $z_{\gamma}\in U_{\gamma}$ and $z_{\delta}\in U_{\delta}$. By 3.11, the first equation in i follows. Since $U_{\beta}$ is abelian, the second equation in i holds as well. Finally, assertion ii follows from i (applied to the $H_{2}$-quintuple $(\epsilon,\delta,\gamma,\beta,\alpha)$) and 3.15:
@@
This is a similar computation as in 5.5. See [Wie24, 7.4.4] for details. ∎
@@
For (ii), recall from 3.10 that $w_{\alpha}$ is also a ($-\alpha$)-Weyl element with factorisation $w_{\alpha}=b_{\alpha}a_{-\alpha}a_{-\alpha}^{w_{\alpha}}$. Hence by two applications of (i),
@@
Using the factorisation $w_{\alpha}^{-1}=(a_{-\alpha}^{-1})^{w_{\alpha}}a_{-\alpha}^{-1}b_{\alpha}^{-1}$ of $w_{\alpha}^{-1}$ as a $(-\alpha)$-Weyl element (see 3.10), an application of (i) to the $H_{2}$-quintuple $(-\alpha,\varepsilon,\delta,\gamma,\beta)$ yields $x_{\varepsilon}=\left[a_{-\alpha}^{-1},[b_{\alpha},x_{\varepsilon}]_{\beta}\right]_{\varepsilon}$. This finishes the proof. ∎
@@
Since $w_{\alpha}$ is also a $(-\alpha)$-Weyl element by 3.10 (iii), we may assume that $\zeta\in\{\beta,\gamma,\delta,\varepsilon\}$. In each case, we will derive a formula for the action of $w_{\alpha}^{2}$ on $U_{\zeta}$ which does not depend on the choice of $w_{\alpha}$, which proves the first assertion. It then follows that $w_{\alpha}^{2}$ and $(w_{\alpha}^{-1})^{2}$ act identically on $U_{\zeta}$, which says precisely that $w_{\alpha}^{4}$ acts trivially on $U_{\zeta}$.
Recall from 2.19 that $\zeta$ lies in a unique $H_{2}$-subsystem $\tilde{\Psi}$ different from $\Psi$. We choose an $H_{2}$-quintuple $(\tilde{\alpha},\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in $\tilde{\Psi}$ such that $\zeta=\tilde{\beta}$. Since any rank-2 subsystem is uniquely determined by two non-proportional roots in this subsystem, we have $\Psi\cap\tilde{\Psi}=\{\zeta,-\zeta\}=\{\tilde{\beta},-\tilde{\beta}\}$ and
@@
Now let $x_{\zeta}\in U_{\zeta}$ and put $w\coloneqq w_{\alpha}^{2}$. By 5.8 we can write $x_{\zeta}=[x_{\tilde{\alpha}},x_{\tilde{\varepsilon}}]_{\zeta}$ for some $x_{\tilde{\alpha}}\in U_{\tilde{\alpha}}$ and $x_{\tilde{\varepsilon}}\in U_{\tilde{\varepsilon}}$. Hence by 3.15,
@@
If $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$, then 5.4 provides formulas for the action of $w$ on $U_{\tilde{\alpha}}$ and $U_{\tilde{\varepsilon}}$ which do not depend on the choice of $w_{\alpha}$. Thus the first assertion holds in this case. Further, since $\alpha$ lies in exactly two $H_{2}$-subsystems by 2.19, it follows from (5.10) that at most one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ can be of type $H_{2}$.
We remain with the case that one of the subsystems $\Psi_{\alpha,\tilde{\alpha}}$ and $\Psi_{\alpha,\tilde{\varepsilon}}$ is of type $H_{2}$. Again by applying 5.8, but to the $H_{2}$-pair $(\tilde{\gamma},-\tilde{\delta})$, we can write $x_{\zeta}=[x_{\tilde{\gamma}},x_{-\tilde{\delta}}]_{\zeta}$ for some $x_{\tilde{\gamma}}\in U_{\tilde{\gamma}}$ and $x_{-\tilde{\delta}}\in U_{-\tilde{\delta}}$. As above, we then have $x_{\zeta}^{w}=[x_{\tilde{\gamma}}^{w},x_{-\tilde{\delta}}^{w}]_{\zeta}$. By our case assumption, 2.19 and (5.10), both $\Psi_{\alpha,\tilde{\gamma}}$ and $\Psi_{\alpha,-\tilde{\delta}}$ are of type $A_{1}\times A_{1}$ or $A_{2}$. By 5.4, we conclude that the first assertion holds in this case as well, which finishes the proof. ∎
@@
By 2.19 iii, there are exactly two positive roots in $H_{3}\cap\alpha^{\perp}$, so the uniqueness assertion is trivial. Let $\rho$ denote any root in $H_{3}\cap\alpha^{\perp}$. Then $\rho\notin\Psi$ and hence $\rho\notin\langle\Psi\rangle_{\mathbb{R}}$. By the defining formula for reflections, it follows that for all $\xi\in\Psi$, we have either $\xi^{\sigma(\rho)}=\xi$ or $\xi^{\sigma(\rho)}\notin\Psi$. At the same time, by 2.19 iii, $\sigma_{\rho}$ cannot fix all roots in $\Psi$. Since $\alpha^{\sigma(\rho)}=\alpha$, we conclude that $\Psi^{\sigma(\rho)}$ is the unique $H_{2}$-subsystem of $H_{3}$ containing $\alpha$ which is different from $\Psi$. That is, $\Psi^{\sigma(\rho)}=\tilde{\Psi}$. Since $\sigma_{\rho}$ is orthogonal and $\alpha^{\sigma(\rho)}=\alpha$, we must have $\beta^{\sigma(\rho)}\in\{\tilde{\beta},-\tilde{\varepsilon}\}$.
@@
Let $\eta,\vartheta\in H_{3}\in\alpha^{\perp}$ be as in 5.14 and choose Weyl elements $w_{\eta}\in M_{\eta}$ and $w_{\vartheta}\in M_{\vartheta}$. At first, we consider the case that $\xi$ and $\zeta$ lie in a common $H_{2}$-subsystem. Without loss of generality, we assume that both lie in $\Psi$. Further, by 3.10 (iii) and 5.9, we may assume that $\zeta\neq\xi$ and that both $\xi,\zeta$ are positive roots. Thus either $\{\xi,\zeta\}=\{\beta,\varepsilon\}$ or $\{\xi,\zeta\}=\{\gamma,\delta\}$. In any case, $\xi^{\sigma(\eta)\sigma(\vartheta)}=-\zeta$ by 5.14, so it follows from 3.10 (ii) that
@@
Note that if $\rho$ has the desired properties, then $-\rho$ has the desired properties for $(\alpha,\tilde{\beta},\tilde{\gamma},\tilde{\delta},\tilde{\varepsilon})$ in 5.13 replaced by $(\alpha,-\tilde{\varepsilon},-\tilde{\delta},-\tilde{\gamma},-\tilde{\beta})$. Thus for each fixed root $\alpha$, there are only four cases to consider: Namely, there are two ways to choose $\Psi$ and two ways to choose an $H_{2}$-quintuple starting from $\alpha$ in $\Psi$. It is a straightforward computation (see 2.18) to prove the assertion in each of these cases for one fixed choice of $\alpha$. Then it follows from 2.14 that it holds for all possible choices of $\alpha$. ∎
@@
Let $x_{\alpha}\in U_{\alpha}$ be arbitrary. Using 5.8, we can write $x_{\alpha}=[x_{\beta},x_{-\gamma}]_{\alpha}$ for some $x_{\beta}\in U_{\beta},~{}x_{-\gamma}\in U_{-\gamma}$. Choose $a_{-\beta}\in U_{-\beta}$ and $b_{\beta}\in U_{\beta}$ such that $w_{\beta}=a_{-\beta}b_{\beta}a_{-\beta}$. Applying 5.7 (ii) to the $(-\beta)$-Weyl element $w_{\beta}=b_{\beta}a_{-\beta}b_{\beta}^{w_{\beta}}$ from 3.10 (iii), we see that
@@
Remark 5.21.
We show how the previous assertions can be extended to $H_{4}$-graded groups $(G^{\prime},(U_{\alpha}^{\prime})_{\alpha\in H_{4}})$. First, we can define the notions of involution position and inverted involution position for roots in $H_{4}$ in the same way as in 5.11. Second, we have to verify that analogues of 5.15 and 5.17 hold. For this, let $\alpha,\xi,\zeta\in H_{4}$ such that $\Psi_{\alpha,\xi}$ and $\Psi_{\alpha,\zeta}$ are of type $H_{2}$ and let $w_{\xi}$, $w_{\zeta}$ be Weyl elements of type $\xi$ and $\zeta$, respectively. If $\Psi_{\alpha,\xi,\zeta}$ is of rank 3, then it must be of type $H_{3}$ because this is the only type of subsystems of $H_{3}$ that contains more than one $H_{2}$-subsystem. If $\Psi_{\alpha,\xi,\zeta}$ is of rank 2, then it is of type $H_{2}$. Hence in any case, $\alpha,\xi,\zeta$ are contained in some $H_{3}$-subsystem of $\Phi$. Thus it follows from 5.15 and 5.17 that $w_{\xi}^{2}$ and $w_{\zeta}^{2}$ act identically on $U_{\alpha}^{X}$ if they are in the same position with respect to $\alpha$, and that their actions differ by inversion if one is in involution position and one is inverted involution position with respect to $\alpha$. This observation allows us to define the twisting involution in $H_{4}$-graded groups as in 5.18. The assertions of 5.19 and 5.20 then remain valid as well.